A country aims to double real GDP per capita in the next 25 years. If the rate of population growth in the country is 1.3% per year then at approximately what rate does real GDP need to grow to achieve this goal?
Answer #1

Answer

by the rule of 70

growth rate to double =70/years=70/25=2.8%

By compounding formula

g=(FV/PV)^(1/n) -1

FV=future value=2

PV=present value=1

n=years=25

g=(2/1)^(1/25) -1

g=0.0281138267=2.81%

=======

real GDP gowth rate =required per capita gowth +population growth rate

=2.8+1.3

=4.1%

the real GDP need to grow to achieve this goal at 4.1%